Circular No. 7022

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998ec IN UGC 3576 Y. L. Qiu, Q. Y. Qiao, and J. Y. Hu, Beijing Astronomical Observatory

(BAO), report their discovery of an apparent supernova in the course of the BAO Supernova Survey with the 0.6-m telescope at Xinglong Station. The new star is located at α = 6^{h}53^{m}06^{s}.11, δ = +50^{o}02′22′′.1 (equinox 2000.0), which is 8′′.7 west and 19′′.5 north of the center of UGC 3576. The object was found and confirmed on unfiltered CCD images taken on Sept. 26.82 and 27.83 UT that show nearly identical magnitudes of 16.9. CCD frames taken on Mar. 22 show no star at the position of the supernova (limiting mag ∼ 18.5).

XTE J1946+274 P. Ghavamian, Rice University; and M. Garcia, Harvard-Smithsonian

Center for Astrophysics, write: “We report photometric and spectrocopic measurements of the source suggested by Israel et al. (IAUC 7021) to be the optical counterpart of the transient x-ray pulsar XTE J1946+274. Spectroscopy (integration time 50 min; spectral resolution 0.6 nm; range 350– 750 nm) was performed on Sept. 27.1 UT with the 1.5-m telescope (+ FAST spectrograph) of the F. L. Whipple Observatory at Mt. Hopkins, using the position reported by Israel et al.. Two point sources appear near that position on the Digital Sky Survey, both of which were covered by the slit (p.a. 90^{o}). The ratio of the Hβ/He I 447.1-nm/Hγ lines indicates a spectral type of B2 (± 2 subclasses) for the brighter of the pair, which is located at α = 19^{h}45^{m}39^{s}.26, δ = +27^{o}22′45′′.3 (equinox 2000.0; ± 0′′.3). Interstellar absorption lines at 577.8, 589.0 (Na D), and 628.4 nm are very prominent, with EW = 0.09, 0.2, and 0.18 nm, respectively. The fainter star lies ∼ 7′′ to the west of the brighter one. The ratio of the Balmer series and Mg b absorption lines in the fainter star suggests a spectral type of F9 ± 1. The FAST spectra of the B star fail to show the weak Hα emission reported by Israel et al.; instead, Hα absorption is seen with EW = 0.68 nm. Photometry was performed on the B star using the CCD imager on the 1.2-m telescope at Mt. Hopkins on Sept. 28.1. Using Guide Star Catalog magnitudes for flux calibration, we find V = 14.4 ± 0.3. The lack of emission lines and variability does not necessarily discount this star as the optical counterpart of XTE J1946+274. The strength of the Na D line indicates that the brighter star is ∼ 4000 pc distant, and implies A(V) ∼ 8. If this star is the optical counterpart of the x-ray pulsar, the x-ray spectrum should show substantial low-energy absorption (N_{H} ∼ 1.5 × 10^{22} cm^{−2}).”

1998 September 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green